OpenStudy (anonymous):

Use matrices to rotate the vector v={-4,1}=by 270°. Write the resulting vector in component form.

OpenStudy (anonymous):

You know the vector rotation matrix?

OpenStudy (anonymous):

I don't :/

OpenStudy (anonymous):

\[\left[\begin{matrix}\cos \theta & -\sin \theta \\ \sin \theta & \cos \theta\end{matrix}\right]\]

OpenStudy (anonymous):

multiply this by {-4,1} where theta is the angle of rotation

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

No not at all, I'm so confused..

OpenStudy (loser66):

@Ready1917 do you know how to multiply 2 matrices?

OpenStudy (anonymous):

noo I don't